You are taking a course that has 4 tests. The first 3 tests are 100 points each and the fourth test is 200 points. To get an A in the course, you must have an average of at least 90% on the four tests. Your scores on the first three tests were 87,92,and 84. What must you score on the 4th test to get an A for the course?
Found 2 solutions by drk, Theo:
Answer by drk(1908)

The total points are 300 + 200 = 500.
90% of 500 = 450 points
---
The 3 scores are: 87, 92,and 84 which sum to 263.
450 - 263 = 187.
---
You must scores 187 out of 200 on the fourth test to get an A.

Answer by Theo(13342)

Your scores are 87, 92, 84.

To get an average of 90 on all four tests, your formula needs to be:

87 + 92 + 84 + x = 90*4

Combine like terms and simplify to get:

263 + x = 360

Subtract 263 from both sides of the equation to get:

x = 360 - 263 = 97.

You need to get a 97 on the last test.

Then you would get the following:


87 + 92 + 84 + 97 = 360

360/4 = 90% overall average on the 4 tests.